# Prototype-Guided Diffusion: Visual Conditioning without External Memory

### v4

@@ VI Conclusion
VII Conclusion and Future Work
We introduced the Prototype Diffusion Model (PDM) and its supervised variant s-PDM, which integrate prototype learning directly into diffusion training. PDM learns prototypes online without labels, offering strong semantic grounding and competitive visual quality while avoiding external retrieval systems and large memory banks. In contrast, s-PDM leverages class labels to attach fixed prototypes, yielding the best quantitative results (e.g., lower FID/KID) but requiring annotations. Our experiments show that prototype-guided conditioning is effective in both regimes: s-PDM delivers state-of-the-art fidelity with labels, while PDM remains a practical, label-free alternative with compelling qualitative results and reduced overhead. Ablations on prototype count further validate that aligning prototype granularity with dataset semantics is key to performance.
Future work will extend prototype-guided conditioning to text-to-image generation by learning a shared multimodal space for language-driven prototype selection and composition. We also plan to explore semi- and weakly supervised variants (bridging PDM and s-PDM), hierarchical/compositional prototypes for complex scenes, and online adaptation to distribution shifts to maintain prototype relevance during long-term training and deployment.

### v5

@@ VI Conclusion
We presented the Prototype Diffusion Model (PDM) and its supervised variant s-PDM, which integrate prototype learning directly into the diffusion process. By jointly updating prototypes with the denoiser, PDM provides label-free semantic guidance without external memory or retrieval, while s-PDM leverages class labels for improved FID/KID performance through stable, class-specific prototypes.
Experiments demonstrate that prototype-guided conditioning consistently enhances generative fidelity: s-PDM achieves the strongest quantitative results, and PDM offers a competitive, annotation-free alternative. Ablations confirm the importance of aligning the number of prototypes with dataset semantics.
Future work includes extending prototype-guided diffusion to text-to-image tasks, exploring semi- and weakly supervised variants, and developing hierarchical or adaptive prototypes for complex and evolving data distributions.